# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ D.1 Results per datasets
Reweighting Dens. RW 0.65 $\pm$ 0.03 0.71 $\pm$ 0.02 0.72 $\pm$ 0.01 0.65 $\pm$ 0.02 0.72 $\pm$ 0.01 0.69 $\pm$ 0.01 0.71 $\pm$ 0.01 0.74 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.71 $\pm$ 0.01 0.7 $\pm$ 0.02 7.05
Disc. RW 0.63 $\pm$ 0.01 0.71 $\pm$ 0.01 0.72 $\pm$ 0.01 0.64 $\pm$ 0.02 0.7 $\pm$ 0.01 0.68 $\pm$ 0.03 0.62 $\pm$ 0.05 0.73 $\pm$ 0.01 0.7 $\pm$ 0.02 0.69 $\pm$ 0.01 0.7 $\pm$ 0.01 0.68 $\pm$ 0.02 7.77
Gauss. RW 0.56 $\pm$ 0.03 0.63 $\pm$ 0.01 0.54 $\pm$ 0.06 0.44 $\pm$ 0.08 0.58 $\pm$ 0.08 0.64 $\pm$ 0.03 0.62 $\pm$ 0.02 0.69 $\pm$ 0.03 0.66 $\pm$ 0.03 0.68 $\pm$ 0.0 0.61 $\pm$ 0.02 0.6 $\pm$ 0.04 16.18
KLIEP 0.65 $\pm$ 0.03 0.69 $\pm$ 0.04 0.72 $\pm$ 0.01 0.66 $\pm$ 0.02 0.72 $\pm$ 0.01 0.69 $\pm$ 0.03 0.68 $\pm$ 0.04 0.75 $\pm$ 0.01 0.65 $\pm$ 0.06 0.68 $\pm$ 0.03 0.71 $\pm$ 0.01 0.69 $\pm$ 0.03 6.64
KMM 0.52 $\pm$ 0.04 0.63 $\pm$ 0.01 0.6 $\pm$ 0.04 0.5 $\pm$ 0.02 0.64 $\pm$ 0.04 0.28 $\pm$ 0.2 0.57 $\pm$ 0.19 0.63 $\pm$ 0.05 0.66 $\pm$ 0.03 0.64 $\pm$ 0.03 0.58 $\pm$ 0.02 0.57 $\pm$ 0.06 17.09
NN RW 0.59 $\pm$ 0.05 0.69 $\pm$ 0.01 0.71 $\pm$ 0.03 0.51 $\pm$ 0.1 0.66 $\pm$ 0.06 0.68 $\pm$ 0.03 0.68 $\pm$ 0.02 0.71 $\pm$ 0.04 0.7 $\pm$ 0.02 0.65 $\pm$ 0.02 0.68 $\pm$ 0.02 0.66 $\pm$ 0.03 10.36
MMDTarS 0.65 $\pm$ 0.03 0.7 $\pm$ 0.02 0.72 $\pm$ 0.01 0.67 $\pm$ 0.01 0.72 $\pm$ 0.01 0.7 $\pm$ 0.01 0.71 $\pm$ 0.02 0.74 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 7.50
Mapping CORAL 0.62 $\pm$ 0.03 0.72 $\pm$ 0.01 0.73 $\pm$ 0.01 0.61 $\pm$ 0.01 0.72 $\pm$ 0.01 0.69 $\pm$ 0.01 0.7 $\pm$ 0.02 0.74 $\pm$ 0.01 0.7 $\pm$ 0.01 0.7 $\pm$ 0.01 0.7 $\pm$ 0.02 0.69 $\pm$ 0.01 4.59
MapOT 0.6 $\pm$ 0.01 0.69 $\pm$ 0.0 0.7 $\pm$ 0.01 0.59 $\pm$ 0.01 0.71 $\pm$ 0.01 0.69 $\pm$ 0.02 0.69 $\pm$ 0.01 0.73 $\pm$ 0.02 0.69 $\pm$ 0.02 0.69 $\pm$ 0.01 0.64 $\pm$ 0.01 0.67 $\pm$ 0.01 9.59
EntOT 0.53 $\pm$ 0.01 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.53 $\pm$ 0.0 0.62 $\pm$ 0.0 0.65 $\pm$ 0.0 0.66 $\pm$ 0.01 0.74 $\pm$ 0.01 0.64 $\pm$ 0.0 0.64 $\pm$ 0.0 0.54 $\pm$ 0.0 0.62 $\pm$ 0.0 14.68
ClassRegOT 0.66 $\pm$ 0.02 0.73 $\pm$ 0.02 0.72 $\pm$ 0.01 0.64 $\pm$ 0.02 0.73 $\pm$ 0.01 0.63 $\pm$ 0.02 0.62 $\pm$ 0.02 0.68 $\pm$ 0.03 0.66 $\pm$ 0.01 0.66 $\pm$ 0.03 0.69 $\pm$ 0.02 0.68 $\pm$ 0.02 12.27
LinOT 0.64 $\pm$ 0.03 0.73 $\pm$ 0.02 0.73 $\pm$ 0.01 0.61 $\pm$ 0.01 0.73 $\pm$ 0.0 0.69 $\pm$ 0.02 0.7 $\pm$ 0.02 0.74 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02 3.59
MMD-LS 0.58 $\pm$ 0.08 0.67 $\pm$ 0.03 0.72 $\pm$ 0.01 0.66 $\pm$ 0.02 0.58 $\pm$ 0.18 0.69 $\pm$ 0.01 0.71 $\pm$ 0.01 0.75 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.71 $\pm$ 0.01 0.68 $\pm$ 0.04 9.14
Subspace JPCA 0.62 $\pm$ 0.02 0.7 $\pm$ 0.01 0.72 $\pm$ 0.01 0.63 $\pm$ 0.01 0.7 $\pm$ 0.02 0.7 $\pm$ 0.01 0.71 $\pm$ 0.01 0.75 $\pm$ 0.01 0.69 $\pm$ 0.02 0.68 $\pm$ 0.01 0.69 $\pm$ 0.03 0.69 $\pm$ 0.02 6.73
SA 0.63 $\pm$ 0.0 0.7 $\pm$ 0.01 0.69 $\pm$ 0.01 0.63 $\pm$ 0.0 0.7 $\pm$ 0.01 0.64 $\pm$ 0.02 0.64 $\pm$ 0.02 0.69 $\pm$ 0.01 0.68 $\pm$ 0.01 0.67 $\pm$ 0.01 0.64 $\pm$ 0.01 0.66 $\pm$ 0.01 10.64
TCA 0.52 $\pm$ 0.0 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.52 $\pm$ 0.0 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.64 $\pm$ 0.0 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.64 $\pm$ 0.0 0.52 $\pm$ 0.0 0.6 $\pm$ 0.0 17.36
TSL 0.65 $\pm$ 0.03 0.71 $\pm$ 0.02 0.72 $\pm$ 0.01 0.66 $\pm$ 0.02 0.72 $\pm$ 0.01 0.69 $\pm$ 0.01 0.71 $\pm$ 0.01 0.75 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.71 $\pm$ 0.01 0.7 $\pm$ 0.01 5.86
Other JDOT 0.61 $\pm$ 0.01 0.7 $\pm$ 0.01 0.71 $\pm$ 0.01 0.6 $\pm$ 0.01 0.7 $\pm$ 0.01 0.63 $\pm$ 0.01 0.64 $\pm$ 0.02 0.72 $\pm$ 0.03 0.69 $\pm$ 0.02 0.69 $\pm$ 0.01 0.67 $\pm$ 0.03 0.67 $\pm$ 0.01 10.27
OTLabelProp 0.59 $\pm$ 0.02 0.69 $\pm$ 0.01 0.7 $\pm$ 0.01 0.59 $\pm$ 0.01 0.7 $\pm$ 0.01 0.69 $\pm$ 0.02 0.68 $\pm$ 0.02 0.74 $\pm$ 0.01 0.69 $\pm$ 0.02 0.69 $\pm$ 0.01 0.63 $\pm$ 0.01 0.67 $\pm$ 0.01 9.41
@@
Table D.6: Accuracy score for Mushrooms dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Reweighting Dens. RW 0.67 $\pm$ 0.01 0.76 $\pm$ 0.0 0.71 $\pm$ 0.01 10.50
Disc. RW 0.73 $\pm$ 0.06 0.78 $\pm$ 0.01 0.75 $\pm$ 0.04 5.00
Gauss. RW 0.56 $\pm$ 0.0 0.46 $\pm$ 0.0 0.51 $\pm$ 0.0 20.75
KLIEP 0.66 $\pm$ 0.02 0.77 $\pm$ 0.01 0.72 $\pm$ 0.01 12.50
KMM 0.7 $\pm$ 0.02 0.78 $\pm$ 0.01 0.74 $\pm$ 0.01 7.00
NN RW 0.67 $\pm$ 0.05 0.75 $\pm$ 0.01 0.71 $\pm$ 0.03 14.00
MMDTarS 0.7 $\pm$ 0.02 0.77 $\pm$ 0.01 0.74 $\pm$ 0.01 7.50
Mapping CORAL 0.66 $\pm$ 0.02 0.77 $\pm$ 0.01 0.72 $\pm$ 0.02 13.50
MapOT 0.65 $\pm$ 0.01 0.62 $\pm$ 0.02 0.63 $\pm$ 0.01 16.50
EntOT 0.82 $\pm$ 0.01 0.67 $\pm$ 0.01 0.75 $\pm$ 0.01 9.00
ClassRegOT 0.92 $\pm$ 0.01 0.72 $\pm$ 0.01 0.82 $\pm$ 0.01 8.00
LinOT 0.72 $\pm$ 0.01 0.81 $\pm$ 0.01 0.76 $\pm$ 0.01 4.50
MMD-LS 0.88 $\pm$ 0.01 0.83 $\pm$ 0.01 0.86 $\pm$ 0.01 3.50
Subspace JPCA 0.76 $\pm$ 0.03 0.81 $\pm$ 0.0 0.78 $\pm$ 0.02 6.00
SA 0.93 $\pm$ 0.05 0.84 $\pm$ 0.01 0.88 $\pm$ 0.03 2.50
TCA 0.45 $\pm$ 0.04 0.46 $\pm$ 0.0 0.45 $\pm$ 0.02 21.75
TSL 0.62 $\pm$ 0.08 0.51 $\pm$ 0.01 0.56 $\pm$ 0.04 19.00
Other JDOT 0.67 $\pm$ 0.02 0.6 $\pm$ 0.01 0.63 $\pm$ 0.01 14.50
OTLabelProp 0.68 $\pm$ 0.01 0.61 $\pm$ 0.01 0.64 $\pm$ 0.01 13.00
DASVM 0.71 $\pm$ 0.02 0.85 $\pm$ 0.0 0.78 $\pm$ 0.01 4.00
Table D.7: Accuracy score for Phishing dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Table D.8: Accuracy score for BCI dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Reweighting Dens. RW 0.59 $\pm$ 0.06 0.51 $\pm$ 0.05 0.72 $\pm$ 0.04 0.5 $\pm$ 0.06 0.39 $\pm$ 0.04 0.4 $\pm$ 0.07 0.63 $\pm$ 0.06 0.67 $\pm$ 0.06 0.53 $\pm$ 0.09 0.55 $\pm$ 0.06 9.72
Disc. RW 0.63 $\pm$ 0.06 0.51 $\pm$ 0.06 0.77 $\pm$ 0.05 0.49 $\pm$ 0.04 0.4 $\pm$ 0.05 0.4 $\pm$ 0.05 0.61 $\pm$ 0.05 0.7 $\pm$ 0.08 0.56 $\pm$ 0.09 0.56 $\pm$ 0.06 6.22
Gauss. RW 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 20.56
KLIEP 0.59 $\pm$ 0.05 0.52 $\pm$ 0.04 0.71 $\pm$ 0.04 0.48 $\pm$ 0.07 0.39 $\pm$ 0.04 0.41 $\pm$ 0.08 0.63 $\pm$ 0.07 0.69 $\pm$ 0.08 0.55 $\pm$ 0.08 0.55 $\pm$ 0.06 8.11
KMM 0.61 $\pm$ 0.06 0.45 $\pm$ 0.08 0.74 $\pm$ 0.05 0.39 $\pm$ 0.1 0.31 $\pm$ 0.09 0.36 $\pm$ 0.05 0.64 $\pm$ 0.09 0.64 $\pm$ 0.01 0.52 $\pm$ 0.07 0.52 $\pm$ 0.06 11.44

### v2

@@ E.1 Results per datasets
Reweighting Dens. RW 0.65 $\pm$ 0.03 0.71 $\pm$ 0.02 0.72 $\pm$ 0.01 0.65 $\pm$ 0.02 0.72 $\pm$ 0.01 0.69 $\pm$ 0.01 0.71 $\pm$ 0.01 0.74 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.71 $\pm$ 0.01 0.7 $\pm$ 0.02 6.82
Disc. RW 0.63 $\pm$ 0.01 0.71 $\pm$ 0.01 0.72 $\pm$ 0.01 0.64 $\pm$ 0.02 0.7 $\pm$ 0.01 0.68 $\pm$ 0.03 0.62 $\pm$ 0.05 0.73 $\pm$ 0.01 0.7 $\pm$ 0.02 0.69 $\pm$ 0.01 0.7 $\pm$ 0.01 0.68 $\pm$ 0.02 7.32
Gauss. RW 0.53 $\pm$ 0.01 0.62 $\pm$ 0.01 0.54 $\pm$ 0.06 0.38 $\pm$ 0.06 0.51 $\pm$ 0.15 NA NA 0.62 $\pm$ 0.01 NA NA NA 0.53 $\pm$ 0.05 19.33
KLIEP 0.65 $\pm$ 0.03 0.69 $\pm$ 0.04 0.72 $\pm$ 0.01 0.66 $\pm$ 0.02 0.72 $\pm$ 0.01 0.69 $\pm$ 0.03 NA 0.75 $\pm$ 0.01 0.65 $\pm$ 0.06 0.68 $\pm$ 0.03 0.71 $\pm$ 0.01 0.69 $\pm$ 0.02 6.30
KMM 0.52 $\pm$ 0.04 0.63 $\pm$ 0.01 0.6 $\pm$ 0.04 0.5 $\pm$ 0.02 0.64 $\pm$ 0.04 0.28 $\pm$ 0.2 NA 0.63 $\pm$ 0.05 0.66 $\pm$ 0.03 0.64 $\pm$ 0.03 0.58 $\pm$ 0.02 0.57 $\pm$ 0.05 16.10
NN RW 0.59 $\pm$ 0.05 0.69 $\pm$ 0.01 0.71 $\pm$ 0.03 0.51 $\pm$ 0.1 0.66 $\pm$ 0.06 0.68 $\pm$ 0.03 NA 0.71 $\pm$ 0.04 0.7 $\pm$ 0.02 0.65 $\pm$ 0.02 0.68 $\pm$ 0.02 0.66 $\pm$ 0.04 10.65
MMDTarS 0.65 $\pm$ 0.03 0.7 $\pm$ 0.02 0.72 $\pm$ 0.01 0.67 $\pm$ 0.01 0.72 $\pm$ 0.01 0.7 $\pm$ 0.01 0.71 $\pm$ 0.02 0.74 $\pm$ 0.01 0.7 $\pm$ 0.03 0.7 $\pm$ 0.02 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 7.55
Mapping CORAL 0.62 $\pm$ 0.03 0.72 $\pm$ 0.01 0.73 $\pm$ 0.01 0.61 $\pm$ 0.01 0.72 $\pm$ 0.01 0.69 $\pm$ 0.01 NA NA 0.7 $\pm$ 0.01 0.7 $\pm$ 0.01 0.7 $\pm$ 0.02 0.69 $\pm$ 0.01 4.56
MapOT 0.6 $\pm$ 0.01 0.7 $\pm$ 0.0 0.71 $\pm$ 0.01 0.59 $\pm$ 0.01 0.7 $\pm$ 0.0 0.68 $\pm$ 0.01 0.69 $\pm$ 0.01 0.73 $\pm$ 0.02 0.69 $\pm$ 0.01 0.69 $\pm$ 0.01 0.64 $\pm$ 0.01 0.67 $\pm$ 0.01 9.36
EntOT 0.53 $\pm$ 0.01 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.53 $\pm$ 0.0 0.62 $\pm$ 0.0 0.65 $\pm$ 0.0 0.66 $\pm$ 0.01 0.74 $\pm$ 0.01 0.64 $\pm$ 0.0 0.64 $\pm$ 0.0 0.54 $\pm$ 0.0 0.62 $\pm$ 0.0 13.86
ClassRegOT 0.68 $\pm$ 0.02 0.72 $\pm$ 0.02 0.69 $\pm$ 0.03 0.67 $\pm$ 0.02 0.72 $\pm$ 0.02 0.61 $\pm$ 0.01 0.61 $\pm$ 0.02 0.69 $\pm$ 0.03 0.65 $\pm$ 0.02 0.65 $\pm$ 0.02 0.7 $\pm$ 0.02 0.67 $\pm$ 0.02 9.00
LinOT 0.64 $\pm$ 0.03 0.73 $\pm$ 0.02 0.73 $\pm$ 0.01 0.61 $\pm$ 0.01 0.73 $\pm$ 0.0 0.69 $\pm$ 0.02 NA 0.74 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02 0.7 $\pm$ 0.02 3.60
MMD-LS 0.56 $\pm$ 0.02 0.54 $\pm$ 0.03 0.72 $\pm$ 0.01 0.63 $\pm$ 0.06 0.61 $\pm$ 0.16 0.65 $\pm$ 0.1 0.55 $\pm$ 0.15 NA 0.63 $\pm$ 0.11 0.48 $\pm$ 0.01 0.71 $\pm$ 0.01 0.61 $\pm$ 0.07 14.30
Subspace JPCA 0.61 $\pm$ 0.0 0.69 $\pm$ 0.01 0.69 $\pm$ 0.01 0.63 $\pm$ 0.01 0.69 $\pm$ 0.0 0.68 $\pm$ 0.01 0.68 $\pm$ 0.01 0.7 $\pm$ 0.01 0.67 $\pm$ 0.01 0.68 $\pm$ 0.01 0.64 $\pm$ 0.01 0.67 $\pm$ 0.01 10.14
SA 0.62 $\pm$ 0.02 0.72 $\pm$ 0.01 0.73 $\pm$ 0.02 0.62 $\pm$ 0.01 0.72 $\pm$ 0.01 0.69 $\pm$ 0.01 0.69 $\pm$ 0.01 0.74 $\pm$ 0.02 0.71 $\pm$ 0.02 0.7 $\pm$ 0.01 0.69 $\pm$ 0.03 0.69 $\pm$ 0.02 8.27
TCA 0.52 $\pm$ 0.0 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.52 $\pm$ 0.0 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.64 $\pm$ 0.0 0.62 $\pm$ 0.0 0.64 $\pm$ 0.0 0.64 $\pm$ 0.0 0.52 $\pm$ 0.0 0.6 $\pm$ 0.0 18.45
TSL 0.65 $\pm$ 0.03 0.71 $\pm$ 0.02 0.72 $\pm$ 0.01 0.66 $\pm$ 0.02 0.72 $\pm$ 0.01 NA NA 0.75 $\pm$ 0.01 0.71 $\pm$ 0.02 0.7 $\pm$ 0.02 0.71 $\pm$ 0.01 0.7 $\pm$ 0.02 5.94
Other JDOT 0.61 $\pm$ 0.01 0.71 $\pm$ 0.01 0.72 $\pm$ 0.01 0.6 $\pm$ 0.01 0.7 $\pm$ 0.0 0.63 $\pm$ 0.01 NA 0.71 $\pm$ 0.02 0.69 $\pm$ 0.02 0.69 $\pm$ 0.01 0.67 $\pm$ 0.03 0.67 $\pm$ 0.01 8.75
OTLabelProp 0.6 $\pm$ 0.01 0.69 $\pm$ 0.01 0.7 $\pm$ 0.01 0.59 $\pm$ 0.01 0.69 $\pm$ 0.01 0.68 $\pm$ 0.02 0.69 $\pm$ 0.02 0.74 $\pm$ 0.01 0.69 $\pm$ 0.01 0.69 $\pm$ 0.0 0.64 $\pm$ 0.01 0.67 $\pm$ 0.01 9.32
@@
Table 13: Accuracy score for Mushrooms dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Reweighting Dens. RW 0.67 $\pm$ 0.01 0.76 $\pm$ 0.0 0.71 $\pm$ 0.01 9.00
Disc. RW 0.73 $\pm$ 0.06 0.78 $\pm$ 0.01 0.75 $\pm$ 0.04 4.50
Gauss. RW 0.56 $\pm$ 0.0 0.46 $\pm$ 0.0 0.51 $\pm$ 0.0 17.50
KLIEP 0.66 $\pm$ 0.02 0.77 $\pm$ 0.01 0.72 $\pm$ 0.01 10.75
KMM 0.7 $\pm$ 0.02 0.78 $\pm$ 0.01 0.74 $\pm$ 0.01 6.00
NN RW 0.67 $\pm$ 0.05 0.75 $\pm$ 0.01 0.71 $\pm$ 0.03 12.00
MMDTarS 0.7 $\pm$ 0.02 0.77 $\pm$ 0.01 0.74 $\pm$ 0.01 6.00
Mapping CORAL 0.66 $\pm$ 0.02 0.77 $\pm$ 0.01 0.72 $\pm$ 0.02 11.50
MapOT 0.65 $\pm$ 0.01 0.62 $\pm$ 0.02 0.63 $\pm$ 0.01 14.00
EntOT 0.82 $\pm$ 0.01 0.67 $\pm$ 0.01 0.75 $\pm$ 0.01 8.00
ClassRegOT 0.63 $\pm$ 0.01 0.62 $\pm$ 0.01 0.62 $\pm$ 0.01 15.50
LinOT 0.72 $\pm$ 0.01 0.81 $\pm$ 0.01 0.76 $\pm$ 0.01 4.00
MMD-LS 0.85 $\pm$ 0.01 NA 0.85 $\pm$ 0.01 4.00
Subspace JPCA 0.64 $\pm$ 0.02 0.78 $\pm$ 0.02 0.71 $\pm$ 0.02 10.00
SA 0.53 $\pm$ 0.02 0.86 $\pm$ 0.01 0.7 $\pm$ 0.02 10.00
Other
OTLabelProp 0.68 $\pm$ 0.01 0.61 $\pm$ 0.01 0.64 $\pm$ 0.01 11.50
Table 14: Accuracy score for Phishing dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Table 15: Accuracy score for BCI dataset for each shift compared for all the methods. A white color means the method does not increase the performance compared to Train Src (Train on the source). Green indicates that the performance improved with the DA methods. The darker the color, the more significant the change.
@@
Reweighting Dens. RW 0.59 $\pm$ 0.06 0.51 $\pm$ 0.05 0.72 $\pm$ 0.04 0.5 $\pm$ 0.06 0.39 $\pm$ 0.04 0.4 $\pm$ 0.07 0.63 $\pm$ 0.06 0.67 $\pm$ 0.06 0.53 $\pm$ 0.09 0.55 $\pm$ 0.06 9.61
Disc. RW 0.63 $\pm$ 0.06 0.51 $\pm$ 0.06 0.76 $\pm$ 0.05 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.42 $\pm$ 0.08 0.6 $\pm$ 0.06 0.68 $\pm$ 0.03 0.56 $\pm$ 0.09 0.56 $\pm$ 0.06 6.89
Gauss. RW 0.28 $\pm$ 0.07 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 0.25 $\pm$ 0.01 20.44
KLIEP 0.59 $\pm$ 0.05 0.52 $\pm$ 0.04 0.71 $\pm$ 0.04 0.48 $\pm$ 0.07 0.39 $\pm$ 0.04 0.41 $\pm$ 0.08 0.63 $\pm$ 0.07 0.69 $\pm$ 0.08 0.55 $\pm$ 0.08 0.55 $\pm$ 0.06 8.17
KMM 0.61 $\pm$ 0.06 0.45 $\pm$ 0.08 0.74 $\pm$ 0.05 0.44 $\pm$ 0.04 0.31 $\pm$ 0.09 0.4 $\pm$ 0.07 0.64 $\pm$ 0.09 0.64 $\pm$ 0.01 0.52 $\pm$ 0.07 0.53 $\pm$ 0.06 10.89